If the middle points of the sides of a triangle are (0,0),(1,2) and (-3,4) then find the area of triangle ?

Grade:11

1 Answers

Arun
6 years ago
As we have read in earlier classes that
area of traingle = 4 (area of triangle formed by joining mid points of sides )
for area of triangle formed by joining mid points of sides – 
as we see the length of all the three sides
lengths are 201/2 , 51/2, 5
hence we see that this forms an right angled triangle
hence area = ½ * 201/2 * 51/2
 = ½ * 1001/2 = ½ * 10 = 5
thus area of triangle = 4 * 5 = 20